# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v4

@@ Appendix E Experimental Setting
We introduce the implementation details in this part. The noise was added to both the training stage and the inference stage. Model details are shown in Table 8 and 9. The image resolution is $224\times 224$ for all the experiments. Pre-trained models on ImageNet-21K are used as the backbone. We train all ResNet and ViT-based models using AdamW optimizer Loshchilov & Hutter (2017). We set the learning rate of each parameter group using a cosine annealing schedule with a minimum of $71e-$ 7. The data augmentation for training only includes the random resized crop and normalization.
@@ 3 $\times$ 3, max pool, stride 2
[…]
@@
[…]
@@ 1 $\times$ 1 average pool, 1000-d fc, softmax
ViT-L 24 1024 4096 16 307M
CNN(ResNet) Setting The training epoch is set to 100. We initialized the learning rate as 00 and linearly increase it to $0.001$ after 10 warmup steps. All the experiments of CNNs are trained on a single Tesla V100 GPU with 32 GB. The batch size for ResNet18, ResNet34, ResNet50, and ResNet101 are 1024, 512, 256, and 128, respectively.
Appendix F More Experiment Results
We present additional experimental results demonstrating the effects of injecting positive noise into various ViT series variants, including SwinTransformer, DeiT, ConViT, and BeiT, using the Tiny ImageNet Le & Yang (2015b) dataset. The positive noise intensity is set to 0.3, and the noise is applied to the final layer. Additionally, we report the performance of NoisyNN across multiple datasets, including Tiny ImageNet Le & Yang (2015b), ImageNet-A Hendrycks et al. (2021), ImageNet-C Hendrycks & Dietterich (2019), CIFAR Krizhevsky (2009), and INbreast Moreira et al. (2012). Furthermore, we compare the proposed method with common data augmentation techniques, alternative noise types, and Manifold MixUp Verma et al. (2019). Lastly, we showcase the application of our method to tasks such as domain generalization and text classification, highlighting its versatility and effectiveness.
F.1 Inject Positive Noise to Variants of ViT
As demonstrated in the paper, the positive noise can be injected into the ViT family. Therefore, in this section, we explore the influence of positive noise on the variants of the ViT. The positive noise used here is identical to that in the paper. For this, we comprehensively compare noise injection to ConViT d’Ascoli et al. (2021), BeiT Bao et al. (2021), DeiT Touvron et al. (2021), and Swin Transformer Liu et al. (2021), and comparisons results are reported in Tabel 10. As expected, these variants of ViTs get benefit from the positive noise. The additional four ViT variants are at the base scale, whose parameters are listed in the table’s last row. For a fair comparison, we use identical experimental settings for each kind of experiment. For example, we use the identical setting for vanilla ConViT, ConViT with different kinds of noise. From the experimental results, we can observe that the different variants of ViT benefit from positive noise and significantly improve prediction accuracy. The results on different scale datasets and variants of the ViT family demonstrate that positive noise can universally improve the model performance by a wide margin.
@@ F.2 Positive Noise on TinyImageNet
Table 12: ViT with different kinds of noise on TinyImageNet. Vanilla means the vanilla model without injecting noise. Accuracy is shown in percentage. Gaussian noise used here is subjected to standard normal distribution. Linear transform noise used in this table is designed to be positive noise. The difference is shown in the bracket. Note ViT-L is overfitting on TinyImageNet Dosovitskiy et al. (2020) Steiner et al. (2021).
@@
F.3 Positive Noise on ImageNet-A
Table 13 shows additional results on ImageNet-A. We further tested the positive linear transformation noise on ImageNet-A, which exhibits a significant domain shift compared to the validation set of ImageNet-1k. The results demonstrate the robustness of our method to domain shift. We also calculate the confusion matrices of our method and ViT-B on ImageNet-A, which are presented in Fig. 3(a) and 3(b), respectively.
Table 13: Top 1 accuracy on ImageNet-A with positive linear transform noise.
Model Top1 Acc. Params. Image Res. Pretrained Dataset
ViT-B 27.4 86M 224 $\times$ 224 ImageNet 21k
NoisyNN (ViT-B based) 34.1 86M 224 $\times$ 224 ImageNet 21k
NoisyNN (ViT-B based) 38.3 348M 384 $\times$ 384 ImageNet 21k
(a) Confusion Matrix of NoisyNN on ImageNet-A.
(b) Confusion Matrix of ViT on ImageNet-A.
Figure 3: Comparison of Confusion Matrices: NoisyNN (ViT-based) and ViT on ImageNet-A.
F.4 Positive Noise on ImageNet-C
Table 14 shows additional results on ImageNet-C. ImageNet-C exhibits various forms of domain shift in comparison to the validation set of ImageNet-1k. The results further demonstrate the robustness of our method to such domain shifts.
Table 14: Top 1 accuracy on ImageNet-C with positive linear transform noise.
Model Top1 Acc. Params. Image Res. Pretrained Dataset
ViT-B 53.4 86M 224 $\times$ 224 ImageNet 21k
NoisyNN (ViT-B based) 58.1 86M 224 $\times$ 224 ImageNet 21k
NoisyNN (ViT-B based) 60.5 348M 384 $\times$ 384 ImageNet 21k
F.5 CIFAR and INbreast Results
Table 15: Comparing ViT-B with NoisyNN on CIFAR-10, CIFAR-100 and INbreast.
Model CIFAR-100 CIFAR-10 INbreast
ViT-B 91.5$\pm$0.1 98.6$\pm$0.1 90.6$\pm$0.2
NoisyNN (ViT-B based) 93.7$\pm$0.1 99.4$\pm$0.1 93.5$\pm$0.1
Results on CIFAR-10, CIFAR-100, and INbreast are shown in Table 15. Showing the effectiveness of NoisyNN beyond ImageNet-based datasets.
F.6 Comparison and Combination with Common Data Augmentation Techniques
We compare our method with common data augmentation methods, and the results are presented in Table 17. Additionally, we combine our method with data augmentations, and the corresponding results are shown in Table 16.
Table 16: Combining NoisyNN with Data Augmentation.
Method ImageNet
NoisyNN (No DA) 89.9$\pm$0.5
NoisyNN + RandomResizedCrop 89.1$\pm$0.5
NoisyNN + RandomHorizontalFlip+RandomResizedCrop 89.2$\pm$0.6
NoisyNN + RandomResizedCrop+RandAugment 89.4$\pm$0.5
Table 17: Comparing NoisyNN with Data Augmentation.
Method ImageNet
ViT-B 84.3
ViT-B+RandomFlip+Gaussian Blur 84.2
ViT-B+RandAugment 85.1
ViT-B+Linear Transformation Noise (NoisyNN) 89.9
F.7 Comparison with Other Noises
Below in Table 18 we compare NoisyNN to other commonly seen noises including White Noise, Uniform Noise and Dropout (Srivastava et al., 2014) on TinyImageNet.
Table 18: Comparison of NoisyNN with other noises on TinyImageNet.
ResNet18 ResNet34 ResNet50
Vanilla 64.01 67.04 69.47
White Noise 64.05 65.97 68.87
Uniform Noise 64.05 66.01 69.01
Gaussian Noise 63.23 64.71 68.17
Salt-and-pepper 55.97 63.52 49.42
Dropout 63.96 67.01 69.40
NoisyNN (ours) 73.32 76.70 76.88
F.8 Positive Noise for Domain Generalization
Domain Generalization (DG) methods try to learn a robust model by training on multiple source domains Volpi et al. (2018); Seo et al. (2020); Carlucci et al. (2019); Huang et al. (2020), while DG methods cannot access the target domains during the training stage. To verify our method in the application of DG tasks, we further conduct experiments on VLCS and PACS, two commonly used datasets in the field of DG. The results are reported in Table 19. As shown in the table, compared to competitive methods, our proposed method achieves state-of-the-art (SOTA) results on the PACS and VLCS datasets.
Table 19: Comparison with other methods in domain generalization tasks.
Method PACS VLCS
ViT Dosovitskiy et al. (2020) (ICLR’21) 85.0 76.9
SDViT (Sultana et al., 2022) (ACCV’22) 88.9 81.9
ALOFT (Guo et al., 2023) (CVPR’23) 91.6 81.3
NoisyViT 93.1 84.4
F.9 Positive Noise for Text Classification
Text classification involves categorizing text into predefined classes or labels (Kowsari et al., 2019). It is widely used in various applications such as spam detection, sentiment analysis, topic labeling, and document categorization. To check whether our method can be applied to a different data modality but within the same problem of classification, we conduct experiments on two popular text classification datasets with widely used models. The results are shown in Table 20. Equipped with our method, TextCNN and TextRNN show a significant improvement in performance.
Table 20: Comparison with other methods in text classification tasks.
Method THUNews AGNews
TextCNN (Kim, 2014) (EMNLP’14) 90.8 89.2
NoisyTextCNN 93.4 89.3
TextRNN (Liu et al., 2016) (IJCAI’16) 90.7 87.7
NoisyTextRNN 95.5 88.1
F.10 Positive Noise for Domain Adaptation
Unsupervised domain adaptation (UDA) aims to learn transferable knowledge across the source and target domains with different distributions Pan & Yang (2009) Wei et al. (2018). There are mainly two kinds of deep neural networks for UDA, which are CNN-based and Transformer-based methods Sun et al. (2022) Yang et al. (2023a). Various techniques for UDA are adopted on these backbone architectures. For example, the discrepancy techniques measure the distribution divergence between source and target domains Long et al. (2018) Sun & Saenko (2016). Adversarial adaptation discriminates domain-invariant and domain-specific representations by playing an adversarial game between the feature extractor and a domain discriminator Ganin & Lempitsky (2015).
Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home Venkateswara et al. (2017) and VisDA2017 Peng et al. (2017). Office-HomeVenkateswara et al. (2017) has 15,500 images of 65 classes from four domains: Artistic (Ar), Clip Art (Cl), Product (Pr), and Real-world (Rw) images. VisDA2017 is a Synthetic-to-Real object recognition dataset, with more than 0.2 million images in 12 classes. We use the ViT-B with a $16\times 16$ patch size, pre-trained on ImageNet. We use minibatch Stochastic Gradient Descent (SGD) optimizer Ruder (2016) with a momentum of 0.9 as the optimizer. The batch size is set to 32. We initialized the learning rate as 0 and linearly warm up to 0.05 after 500 training steps. The results are shown in Table 21 and 22. The methods above the black line are based on CNN architecture, while those under the black line are developed from the Transformer architecture. The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results show that positive noise can improve model generality and, therefore, benefit deep models in domain adaptation tasks.
Table 21: Comparison with SOTA methods (ResNet-50 He et al. (2016), MinEnt Grandvalet & Bengio (2004), SAFN Xu et al. (2019), CDAN+E Long et al. (2018), DCAN Li et al. (2020), BNM Cui et al. (2020), SHOT Liang et al. (2020), ATDOC-NA Liang et al. (2021), ViT-B Dosovitskiy et al. (2020), TVT-B Yang et al. (2023a), CDTrans-B Xu et al. (2022), SSRT-B Sun et al. (2022)) on Office-Home. The best performance is marked in red.
@@ F.10 Positive Noise for Domain Adaptation
ResNet-50 44.9 66.3 74.3 51.8 61.9 63.6 52.4 39.1 71.2 63.8 45.9 77.2 59.4
MinEnt 51.0 71.9 77.1 61.2 69.1 70.1 59.3 48.7 77.0 70.4 53.0 81.0 65.8
SAFN 52.0 71.7 76.3 64.2 69.9 71.9 63.7 51.4 77.1 70.9 57.1 81.5 67.3
CDAN+E 54.6 74.1 78.1 63.0 72.2 74.1 61.6 52.3 79.1 72.3 57.3 82.8 68.5
DCAN 54.5 75.7 81.2 67.4 74.0 76.3 67.4 52.7 80.6 74.1 59.1 83.5 70.5
BNM 56.7 77.5 81.0 67.3 76.3 77.1 65.3 55.1 82.0 73.6 57.0 84.3 71.1
SHOT 57.1 78.1 81.5 68.0 78.2 78.1 67.4 54.9 82.2 73.3 58.8 84.3 71.8
ATDOC-NA 58.3 78.8 82.3 69.4 78.2 78.2 67.1 56.0 82.7 72.0 58.2 85.5 72.2
ViT-B 54.7 83.0 87.2 77.3 83.4 85.6 74.4 50.9 87.2 79.6 54.8 88.8 75.5
TVT-B 74.9 86.8 89.5 82.8 88.0 88.3 79.8 71.9 90.1 85.5 74.6 90.6 83.6
CDTrans-B 68.8 85.0 86.9 81.5 87.1 87.3 79.6 63.3 88.2 82.0 66.0 90.6 80.5
SSRT-B 75.2 89.0 91.1 85.1 88.3 90.0 85.0 74.2 91.3 85.7 78.6 91.8 85.4
@@
ResNet-50He et al. (2016) 55.1 53.3 61.9 59.1 80.6 17.9 79.7 31.2 81.0 26.5 73.5 8.5 52.4
DANNGanin & Lempitsky (2015) 81.9 77.7 82.8 44.3 81.2 29.5 65.1 28.6 51.9 54.6 82.8 7.8 57.4
MinEntGrandvalet & Bengio (2004) 80.3 75.5 75.8 48.3 77.9 27.3 69.7 40.2 46.5 46.6 79.3 16.0 57.0
SAFNXu et al. (2019) 93.6 61.3 84.1 70.6 94.1 79.0 91.8 79.6 89.9 55.6 89.0 24.4 76.1
CDAN+ELong et al. (2018) 85.2 66.9 83.0 50.8 84.2 74.9 88.1 74.5 83.4 76.0 81.9 38.0 73.9
BNM Cui et al. (2020) 89.6 61.5 76.9 55.0 89.3 69.1 81.3 65.5 90.0 47.3 89.1 30.1 70.4
CGDMDu et al. (2021) 93.7 82.7 73.2 68.4 92.9 94.5 88.7 82.1 93.4 82.5 86.8 49.2 82.3
SHOTLiang et al. (2020) 94.3 88.5 80.1 57.3 93.1 93.1 80.7 80.3 91.5 89.1 86.3 58.2 82.9
ViT-BDosovitskiy et al. (2020) 97.7 48.1 86.6 61.6 78.1 63.4 94.7 10.3 87.7 47.7 94.4 35.5 67.1
TVT-BYang et al. (2023a) 92.9 85.6 77.5 60.5 93.6 98.2 89.4 76.4 93.6 92.0 91.7 55.7 83.9
CDTrans-BXu et al. (2022) 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
SSRT-B Sun et al. (2022) 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8

### v5

@@ Appendix E Experimental Setting
We introduce the implementation details in this part. There are three different noise injection modes, as shown in Fig. 3. We follow Mode 3 in this paper. Model details are shown in Table 8 and 9. The image resolution is $224\times 224$ for all the experiments. Pre-trained models on ImageNet-21K are used as the backbone. We train all ResNet and ViT-based models using AdamW optimizer [35]. We set the learning rate of each parameter group using a cosine annealing schedule with a minimum of 1e71e-7. The data augmentation for training only includes the random resized crop and normalization.
@@ 3 $\times$ 3, max pool, stride 2
$\begin{bmatrix}1\times 1&128\\ \ 3\times 3&128\\ 1\times 1&512\end{bmatrix}\times 4$
$\begin{bmatrix}1\times 1&128\\ \ 3\times 3&128\\ 1\times 1&512\end{bmatrix}\times 4$
@@
$\begin{bmatrix}1\times 1&256\\ \ 3\times 3&256\\ 1\times 1&1024\end{bmatrix}\times 23$
@@ 1 $\times$ 1 average pool, 1000-d fc, softmax
ViT-L 12 1024 4096 16 307M
@@
Table 12: ViT with different kinds of noise on TinyImageNet. Vanilla means the vanilla model without injecting noise. Accuracy is shown in percentage. Gaussian noise used here is subjected to standard normal distribution. Linear transform noise used in this table is designed to be positive noise. The difference is shown in the bracket. Note ViT-L is overfitting on TinyImageNet [15] [54].
@@
CNN(ResNet) Setting The training epoch is set to 100. We initialized the learning rate as 0 and linearly increase it to $0.001$ after 10 warmup steps. All the experiments of CNNs are trained on a single Tesla V100 GPU with 32 GB. The batch size for ResNet18, ResNet34, ResNet50, and ResNet101 are 1024, 512, 256, and 128, respectively.
Appendix F More Experiment Results
We show more experiment results of injecting positive noise to other variants of the ViT family, such as SwinTransformer, DeiT, ConViT, and BeiT, and implement them on the smaller dataset, i.e., TinyImageNet. Note, considering limited computational resources, all the experiments in the supplementary are on the TinyImageNet. The strength of positive noise is set to 0.3. The noise is injected into the last layer.
F.1 Inject Positive Noise to Variants of ViT
As demonstrated in the paper, the positive noise can be injected into the ViT family. Therefore, in this section, we explore the influence of positive noise on the variants of the ViT. The positive noise used here is identical to that in the paper. For this, we comprehensively compare noise injection to ConViT [11], BeiT [3], DeiT [59], and Swin Transformer [33], and comparisons results are reported in Tabel 10. As expected, these variants of ViTs get benefit from the positive noise. The additional four ViT variants are at the base scale, whose parameters are listed in the table’s last row. For a fair comparison, we use identical experimental settings for each kind of experiment. For example, we use the identical setting for vanilla ConViT, ConViT with different kinds of noise. From the experimental results, we can observe that the different variants of ViT benefit from positive noise and significantly improve prediction accuracy. The results on different scale datasets and variants of the ViT family demonstrate that positive noise can universally improve the model performance by a wide margin.
F.2 Positive Noise on TinyImageNet
Table 13: Comparison with SOTA methods on Office-Home. The best performance is marked in red.
@@ F.2 Positive Noise on TinyImageNet
ResNet-50[22] 44.9 66.3 74.3 51.8 61.9 63.6 52.4 39.1 71.2 63.8 45.9 77.2 59.4
MinEnt[21] 51.0 71.9 77.1 61.2 69.1 70.1 59.3 48.7 77.0 70.4 53.0 81.0 65.8
SAFN[65] 52.0 71.7 76.3 64.2 69.9 71.9 63.7 51.4 77.1 70.9 57.1 81.5 67.3
CDAN+E[34] 54.6 74.1 78.1 63.0 72.2 74.1 61.6 52.3 79.1 72.3 57.3 82.8 68.5
DCAN[29] 54.5 75.7 81.2 67.4 74.0 76.3 67.4 52.7 80.6 74.1 59.1 83.5 70.5
BNM [10] 56.7 77.5 81.0 67.3 76.3 77.1 65.3 55.1 82.0 73.6 57.0 84.3 71.1
SHOT[31] 57.1 78.1 81.5 68.0 78.2 78.1 67.4 54.9 82.2 73.3 58.8 84.3 71.8
ATDOC-NA[32] 58.3 78.8 82.3 69.4 78.2 78.2 67.1 56.0 82.7 72.0 58.2 85.5 72.2
ViT-B[15] 54.7 83.0 87.2 77.3 83.4 85.6 74.4 50.9 87.2 79.6 54.8 88.8 75.5
TVT-B[67] 74.9 86.8 89.5 82.8 88.0 88.3 79.8 71.9 90.1 85.5 74.6 90.6 83.6
CDTrans-B[66] 68.8 85.0 86.9 81.5 87.1 87.3 79.6 63.3 88.2 82.0 66.0 90.6 80.5
SSRT-B [57] 75.2 89.0 91.1 85.1 88.3 90.0 85.0 74.2 91.3 85.7 78.6 91.8 85.4
@@
ResNet-50[22] 55.1 53.3 61.9 59.1 80.6 17.9 79.7 31.2 81.0 26.5 73.5 8.5 52.4
DANN[19] 81.9 77.7 82.8 44.3 81.2 29.5 65.1 28.6 51.9 54.6 82.8 7.8 57.4
MinEnt[21] 80.3 75.5 75.8 48.3 77.9 27.3 69.7 40.2 46.5 46.6 79.3 16.0 57.0
SAFN[65] 93.6 61.3 84.1 70.6 94.1 79.0 91.8 79.6 89.9 55.6 89.0 24.4 76.1
CDAN+E[34] 85.2 66.9 83.0 50.8 84.2 74.9 88.1 74.5 83.4 76.0 81.9 38.0 73.9
BNM [10] 89.6 61.5 76.9 55.0 89.3 69.1 81.3 65.5 90.0 47.3 89.1 30.1 70.4
CGDM[16] 93.7 82.7 73.2 68.4 92.9 94.5 88.7 82.1 93.4 82.5 86.8 49.2 82.3
SHOT[31] 94.3 88.5 80.1 57.3 93.1 93.1 80.7 80.3 91.5 89.1 86.3 58.2 82.9
ViT-B[15] 97.7 48.1 86.6 61.6 78.1 63.4 94.7 10.3 87.7 47.7 94.4 35.5 67.1
TVT-B[67] 92.9 85.6 77.5 60.5 93.6 98.2 89.4 76.4 93.6 92.0 91.7 55.7 83.9
CDTrans-B[66] 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
SSRT-B [57] 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8
@@
F.3 Positive Noise for Domain Adaptation
Unsupervised domain adaptation (UDA) aims to learn transferable knowledge across the source and target domains with different distributions [42] [63]. There are mainly two kinds of deep neural networks for UDA, which are CNN-based and Transformer-based methods [57] [67]. Various techniques for UDA are adopted on these backbone architectures. For example, the discrepancy techniques measure the distribution divergence between source and target domains [34] [55]. Adversarial adaptation discriminates domain-invariant and domain-specific representations by playing an adversarial game between the feature extractor and a domain discriminator [19].
Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home [62] and VisDA2017 [43]. Office-Home[62] has 15,500 images of 65 classes from four domains: Artistic (Ar), Clip Art (Cl), Product (Pr), and Real-world (Rw) images. VisDA2017 is a Synthetic-to-Real object recognition dataset, with more than 0.2 million images in 12 classes. We use the ViT-B with a $16\times 16$ patch size, pre-trained on ImageNet. We use minibatch Stochastic Gradient Descent (SGD) optimizer [46] with a momentum of 0.9 as the optimizer. The batch size is set to 32. We initialized the learning rate as 0 and linearly warm up to 0.05 after 500 training steps. The results are shown in Table 13 and 14. The methods above the black line are based on CNN architecture, while those under the black line are developed from the Transformer architecture. The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results show that positive noise can improve model generality and, therefore, benefit deep models in domain adaptation tasks.